A Fiches enseignants-chercheurs

1

A.1 Catuscia Palamidessi

Statut  : DR  Section du CNU : 27  HDR : Oui
Établissement : INRIA et Ecole Polytechnique
Laboratoire : Futurs et Laboratoire d’Informatique de l’X (LIX)
Équipe  : Comète

Thèmes de recherche

– Theory of Concurrency.

– Security.

– Automatic verification (in particular, model checking).

– Probabilistic systems.

Publications les plus significatives

[1] K. Chatzikokolakis, C. Palamidessi, and P. Panangaden. Anonymity protocols as noisy channels. Information and Computation, 2007. To appear. The report version is available at

www.lix.polytechnique.fr/∼catuscia/papers/Anonymity/Channels/full.pdf.

[2] K. Chatzikokolakis, C. Palamidessi, and P. Panangaden. Probability of error in informationhiding protocols. In Proceedings of the 20th IEEE Computer Security Foundations Symposium (CSF20), pages 341–354. IEEE Computer Society, 2007.

[3] C. Palamidessi, V. A. Saraswat, F. D. Valencia, and B. Victor. On the expressiveness of linearity vs persistence in the asynchronous pi-calculus. In Proceedings of the Twenty First Annual IEEE Symposium on Logic in Computer Science (LICS), pages 59–68. IEEE Computer Society, 2006.

[4] C. Palamidessi and O. M. Herescu. A randomized encoding of the π-calculus with mixed choice. Theoretical Computer Science, 335(2-3) :373–404, 2005.

[5] C. Palamidessi. Comparing the expressive power of the synchronous and the asynchronous pi-calculus. Mathematical Structures in Computer Science, 13(5) :685–719, 2003.

Responsabilités pédagogiques, administratives, animation de la recherche

– Director of the INRIA Equipe Comète

– Member of – The Council of the EATCS (European Association of Theoretical Computer Science)

– The IFIP Technical Committee 1 – Foundations of Computer Science.

– The IFIP Working Group 2.2

– Member of the Editorial Board of – Mathematical Structures in Computer Science, Cambridge University Press.

– Theory and Practice of Logic Programming, Cambridge University Press.

– Electronic Notes of Theoretical Computer Science, Elsevier Science.

Doctorants encadrés depuis 2002

– Oltea Michaela Herescu, The Probabilistic Asynchronous π-calculus, soutenance novembre 2002.

– Konstantinos Chatzikokolakis, Probabilistic and Information-Theoretic Approaches to Anonymity, soutenance octobre 2007.

– Romain Beauxis, Asynchronous Processes for Security, début en novembre 2005.

– Sylvain Pradalier, Stochastic Process Calculi and applications to Bioinformatics, début en mai 2006. Thèse co-encadrée avec Cosimo Laneve, University of Bologna.

– Carlos Olarte, Probabilistic Concurrent Constraints and applications to Security, début en septembre 2006. Thèse co-encadrée avec Frank Valencia, LIX.

– Jesus Aranda, Expressiveness in Process Calculi, début en septembre 2006. Thèse coencadrée avec Frank Valencia, LIX and Juan Francisco Diaz, Universidad del Valle, Colombia.

– Christelle Braun, An epistemological framework for Probabilistic Information Flow, début en octobre 2007.